<!DOCTYPE HTML PUBLIC "-//W3C//DTD HTML 4.01 Transitional//EN">
<html>
  <head>

    <meta http-equiv="content-type" content="text/html; charset=ISO-8859-1">
  </head>
  <body>
    <title>B-team Meeting Notes</title>
    <meta http-equiv="Content-Type" content="text/html;
      charset=ISO-8859-1">
    <meta name="generator" content="Org-mode">
    <meta name="generated" content="2010-12-13 15:37:41 EST">
    <meta name="author" content="Arne Freyberger">
    <meta name="description" content="">
    <meta name="keywords" content="">
    <style type="text/css">
 <!--/*--><![CDATA[/*><!--*/
  html { font-family: Times, serif; font-size: 12pt; }
  .title  { text-align: center; }
  .todo   { color: red; }
  .done   { color: green; }
  .tag    { background-color: #add8e6; font-weight:normal }
  .target { }
  .timestamp { color: #bebebe; }
  .timestamp-kwd { color: #5f9ea0; }
  p.verse { margin-left: 3% }
  pre {
        border: 1pt solid #AEBDCC;
        background-color: #F3F5F7;
        padding: 5pt;
        font-family: courier, monospace;
        font-size: 90%;
        overflow:auto;
  }
  table { border-collapse: collapse; }
  td, th { vertical-align: top; }
  dt { font-weight: bold; }
  div.figure { padding: 0.5em; }
  div.figure p { text-align: center; }
  .linenr { font-size:smaller }
  .code-highlighted {background-color:#ffff00;}
  .org-info-js_info-navigation { border-style:none; }
  #org-info-js_console-label { font-size:10px; font-weight:bold;
                               white-space:nowrap; }
  .org-info-js_search-highlight {background-color:#ffff00; color:#000000;
                                 font-weight:bold; }
  /*]]>*/-->
</style>
    <script type="text/javascript">
<!--/*--><![CDATA[/*><!--*/
 function CodeHighlightOn(elem, id)
 {
   var target = document.getElementById(id);
   if(null != target) {
     elem.cacheClassElem = elem.className;
     elem.cacheClassTarget = target.className;
     target.className = "code-highlighted";
     elem.className   = "code-highlighted";
   }
 }
 function CodeHighlightOff(elem, id)
 {
   var target = document.getElementById(id);
   if(elem.cacheClassElem)
     elem.className = elem.cacheClassElem;
   if(elem.cacheClassTarget)
     target.className = elem.cacheClassTarget;
 }
/*]]>*///-->
</script>
    <div id="content">
      <div align="left">
      </div>
      Also attached are my running notes on the Arc1 optics, see final
      entry of these notes for some late breaking news.  The prioritized
      BSY diagnostics package for FY11 is also attached.<br>
      Arne<br>
      <h1 class="title">B-team Meeting Notes</h1>
      <div id="outline-container-1" class="outline-2">
        <h2 id="sec-1"><span class="section-number-2">1</span> <span
            class="timestamp-wrapper"> <span class="timestamp">2010-12-08
              Wed 14:00</span></span> BTeam Meeting–Notes </h2>
        <div class="outline-text-2" id="text-1">
        </div>
        <div id="outline-container-1.1" class="outline-3">
          <h3 id="sec-1.1"><span class="section-number-3">1.1</span>
            Optics-on-Call Report    <span class="tag"><span
                class="Kazimi">Kazimi</span> <span class="Hofler">Hofler</span></span></h3>
          <div class="outline-text-3" id="text-1.1">
            <p>No calls received. Support Injector transport after the
              cessiate-bake effort on Tuesday.
              150μ A setup complete with more interception than previous
              setup.
            </p>
          </div>
        </div>
        <div id="outline-container-1.2" class="outline-3">
          <h3 id="sec-1.2"><span class="section-number-3">1.2</span>
            Beam Studies Report </h3>
          <div class="outline-text-3" id="text-1.2">
            <ul>
              <li id="sec-1.2.1">Summary of Heat-Activate <br>
                Pockel Cell alignment did not take place. Plenty of
                plenary steps,
                but no alignment.
                <p>
                  Re-thinking the requirements for this phase of the
                  QWeak experiment.
                  Would like a few days of the Winter break. EGG group
                  will support
                  this. What else do we need? </p>
                <p>
                  Next heat-cessiate party will be during the Winter
                  Break.
                </p>
              </li>
              <li id="sec-1.2.2">Parity Quality Beam Status Report    <span
                  class="tag"><span class="Qweak">Qweak</span></span><br>
                Charge asymmetry data. Short periods where the A<sub>Q</sub>
                is not controlled
                due to data acquisition issues. But when the feedback
                loop is working
                A<sub>Q</sub> is reasonably stable.
                <p>
                  Position differences are small when the A<sub>Q</sub>
                  is stable. Future
                  position differences will be for 1h slugs of data.
                  Expect the
                  differences to be in the 10s of nm. Present position
                  differences
                  good enough for this phase of the experiment.
                </p>
                <p>
                  Present asymmetries look encouraging for the run out
                  through May
                  13<sup>th</sup>
                </p>
                <ul>
                  <li id="sec-1.2.2.1">Compton Signal <br>
                    Signal to background looks very good except that
                    there is a 20minute
                    rise and fall of the background. Not sure where the
                    background comes
                    from.
                    <p>
                      Speculation is that it is correlated with the QE
                      hole on the cathode.
                    </p>
                  </li>
                  <li id="sec-1.2.2.2">Things that still need to be
                    checked out <br>
                    <ul>
                      <li>
                        modulation system and hall optics properly
                        configure
                      </li>
                      <li>
                        BCM noise issue (EES involved)
                      </li>
                      <li>
                        Measure betatron damping via PZT
                      </li>
                      <li>
                        optimize FFB FF system, what are the
                        requirements? If the
                        asymmetries look ok do we really need to improve
                        the FFB FF system?
                        "I would not mess with it", says Roger Carlini.
                      </li>
                      <li>
                        IA feedback
                      </li>
                      <li>
                        0.2 Hz kick? Still present in the Injector. Hall
                        needs to
                        evaluate the impact of this and report back.
                        This has been a
                        nagging issue since PRex. </li>
                    </ul>
                  </li>
                </ul>
              </li>
            </ul>
          </div>
        </div>
        <div id="outline-container-1.3" class="outline-3">
          <h3 id="sec-1.3"><span class="section-number-3">1.3</span>
            Beam Studies Planning </h3>
          <div class="outline-text-3" id="text-1.3">
            <ul>
              <li>
                RF Recovery
              </li>
              <li>
                High Current (180 {μ}A) Injector setup for QWeak
              </li>
              <li>
                Mott Measurements
              </li>
              <li>
                Arc9 Transport
              </li>
            </ul>
          </div>
        </div>
        <div id="outline-container-1.4" class="outline-3">
          <h3 id="sec-1.4"><span class="section-number-3">1.4</span> BSY
            Diagnostics    <span class="tag"><span class="Benesch">Benesch</span></span></h3>
          <div class="outline-text-3" id="text-1.4">
            <p>Present updated table. B-team to prioritize the list
              based on
              desirability and known resource limits. The outcome of
              this
              discussion will be the plan that is executed.
            </p>
            <p>
              See Jay's attached list with the prioritized list.
            </p>
            <ul>
              <li id="sec-1.4.1"><span class="todo TODO"> TODO</span>
                Need to finish ItoV on the dumplettes <br>
              </li>
            </ul>
          </div>
        </div>
        <div id="outline-container-1.5" class="outline-3">
          <h3 id="sec-1.5"><span class="section-number-3">1.5</span>
            12GeV Pathlength Requirements Status    <span class="tag"><span
                class="Tiefenback">Tiefenback</span></span></h3>
          <div class="outline-text-3" id="text-1.5">
            <p>Update on requirements table and overall documentation
              preparation.
            </p>
          </div>
        </div>
        <div id="outline-container-1.6" class="outline-3">
          <h3 id="sec-1.6"><span class="section-number-3">1.6</span>
            ARC1 model    <span class="tag"><span class="Arne">Arne</span></span></h3>
          <div class="outline-text-3" id="text-1.6">
            <p>Results of the latest round of Arc1 modeling. See
              separate note. Recent developments have identified the
              problem
              with getting the Arc1 transport to be symmetric. The body
              gradient
              for MBE1A01 is unique when compared to the other 15dipoles
              in the arc
              which all have the same body gradient. This problem is in
              the Art++
              decks as well. See separate note for details.
            </p>
            <p>
              The 12 GeV decks have used the MBE1A01 value for all 16
              dipoles in the
              Arc, whereas a better choice might be to set MBE1A01 to
              the values
              found for the other 15 dipoles. Waiting guidance on this….
            </p>
          </div>
        </div>
      </div>
      <div id="outline-container-2" class="outline-2">
        <h2 id="sec-2"><span class="section-number-2">2</span> <span
            class="timestamp-wrapper"> <span class="timestamp">2010-12-01
              Wed 14:00</span></span> BTeam Meeting–Notes </h2>
        <div class="outline-text-2" id="text-2">
        </div>
        <div id="outline-container-2.1" class="outline-3">
          <h3 id="sec-2.1"><span class="section-number-3">2.1</span>
            Optics-on-Call Report: Hofler </h3>
          <div class="outline-text-3" id="text-2.1">
            <p>Summary of the thankful week
            </p>
            <p>
              Fri: Hall C Compton scan – done.
            </p>
            <p>
              Weekend: Energy stability problems, gang phase lock still
              a problem? A power cycle may be required; readback may
              be invalid; also an observed current dependence in
              tracking.
            </p>
            <p>
              Wed: Hall C Compton backed out, halo counts up, fixed by
              rephasing. Need to get back through Compton. Still an open
              issue, should revisit after spot move on Thursday 12-02.
            </p>
          </div>
        </div>
        <div id="outline-container-2.2" class="outline-3">
          <h3 id="sec-2.2"><span class="section-number-3">2.2</span>
            Beam Studies Report: All </h3>
          <div class="outline-text-3" id="text-2.2">
            <ul>
              <li id="sec-2.2.1">Pathlength Analysis: Yunn <br>
                Status of the analysis effort of the tunnel
                heating/pathlength
                five-pass test in September, final report due next week.
                Byung's
                draft report is attached. Fig 1 shows slower cooling
                than warming.
                12 GeV conditions will likely give RF shift of 2-3
                degrees/arc.
                15F LCW rise was half of 12 GeV conditions; concrete
                temp did not
                asymptote. Should we reperform this study with a longer
                test and/or
                higher temperature? Is this enough data to draw
                conclusions?
                <p>
                  Should we get a baseline to separate out baseline
                  pathlength
                  changes vs temperature-driven changes? This is small
                  compared
                  to expected effect. Estimate that these are acceptable
                  but more
                  numbers on air/wall temp and air conditioning
                  required. Should
                  pull sensors for data and compare, and plan to deliver
                  conclusions
                  before winter break (Tief).
                </p>
              </li>
              <li id="sec-2.2.2">Arc9 model calibration test: Roblin <br>
                <a
                  href="http://opsweb.acc.jlab.org/CSUEApps/atlis/task/10511">ATLiS
                  entry 10511</a>. Multiple RF FSD trips, ran fopt three
                times, not
                necessarily complete, requires data quality review by
                Yves.
              </li>
              <li id="sec-2.2.3">Hall C double peak source
                found/resolved? Suleiman <br>
                <a
href="http://opweb.acc.jlab.org/CSUEApps/elog/entry/1555706?PHPSESSID=eba0eb97d22d5f6c7ee0b57e9c387545">See
                  this elog entry</a>. Can we get a
                troubleshooting/diagnostic?
                Or have we lost some troubleshooting for low current by
                turning
                off the beam scraping monitor? It is a useful archived
                diagnostic
                and should evaluate whether we can turn it back on in
                future.
                We should ensure that it stays off until we want it back
                on.
                Was also used as low-level transfer function monitor,
                monitoring
                for changes in lattice. We will return to this in the
                future to
                see if it can be acceptably turned back on. We'll modify
                app and
                send signals to Qweak for monitoring in the meantime.
              </li>
            </ul>
          </div>
        </div>
        <div id="outline-container-2.3" class="outline-3">
          <h3 id="sec-2.3"><span class="section-number-3">2.3</span>
            Beam Studies Planning: All </h3>
          <div class="outline-text-3" id="text-2.3">
            <ul>
              <li id="sec-2.3.1">Thursday 2010-12-02 and Tuesday
                2010-12-07 Plans: Suggested <br>
                <ul>
                  <li id="sec-2.3.1.1">Tuesday 2010-12-07 <br>
                    CHL power switch uncertainties, injector cathode
                    activation (~12h).
                    TPE also scheduled to start operations at start of
                    evening shift.
                  </li>
                  <li id="sec-2.3.1.2">Compton backgrounds (hall A) <br>
                  </li>
                  <li id="sec-2.3.1.3">Injector development for Hall
                    C/Qweak (Dalton, Benesch), ~4h (to 12h?) <br>
                    <a
                      href="http://opsweb.acc.jlab.org/CSUEApps/atlis/task/10271">ATLiS
                      entry 10271</a>, Rebalance systematics, RHWP
                    scans/PC translation scans.
                    Was done before but with IHWP in wrong orientation.
                    Oops! May require
                    some time both days.
                    <p>
                      Estimate 4h Thu 12-02, 8h Tue 12-07 (laser table
                      in conjunction with
                      Cathode activation, total time 12h)
                    </p>
                  </li>
                  <li id="sec-2.3.1.4">One pass stripline BPM test, ~1h
                    <br>
                    <a
                      href="http://opsweb.acc.jlab.org/CSUEApps/atlis/task/10445">ATLiS
                      entry 10445</a>, Large dynamic range of currents
                    (100uA to >=100 nA CW)
                    for BPM linearity test. Access required, not time
                    critical; can wait
                    for next access.
                  </li>
                  <li id="sec-2.3.1.5">Repeat of Arc9 model calibration
                    test?, ~2h <br>
                    <a
                      href="http://opsweb.acc.jlab.org/CSUEApps/atlis/task/10511">ATLiS
                      entry 10511,</a> should wait for Yves evaluation
                    of existing data to see
                    if its enough for his analysis.
                  </li>
                  <li id="sec-2.3.1.6">Cryomodule JT valve test?, ~10h <br>
                    <a
                      href="http://opsweb.acc.jlab.org/CSUEApps/atlis/task/7874">ATLiS
                      entry 7874</a>, Has been lingering since June.
                    Requires beam off.
                    Also requires Jonathan Creel's availability.
                  </li>
                </ul>
              </li>
              <li id="sec-2.3.2">Thursday 2010-12-02 Plans: Decided <br>
                <ul>
                  <li id="sec-2.3.2.1">Move spot (0.5h) <br>
                  </li>
                  <li id="sec-2.3.2.2">Injector development for QWeak
                    (4h) <br>
                  </li>
                  <li id="sec-2.3.2.3">Compton background tuning
                    (0.5-1h) <br>
                  </li>
                </ul>
              </li>
              <li id="sec-2.3.3">Tuesday 2010-12-07 Plans: Decided <br>
                <ul>
                  <li id="sec-2.3.3.1">Injector development for QWeak,
                    laser table, cathode activation (8-12h) <br>
                  </li>
                </ul>
              </li>
            </ul>
          </div>
        </div>
        <div id="outline-container-2.4" class="outline-3">
          <h3 id="sec-2.4"><span class="section-number-3">2.4</span>
            FY11 BTeam Projects </h3>
          <div class="outline-text-3" id="text-2.4">
            <ul>
              <li id="sec-2.4.1">BSY Diagnostics: Benesch <br>
                Details of the FY11 BSY diagnostics effort: showed list
                of 5 diagnostics
                suggestions on overhead (FY11 diagnostic suggestions).
                Discussions of
                whether to relocate harps back to E03 crosses, or
                install new crosses
                in front of S10 girders for optics matching after each
                recombiner for
                12 GeV upgrade commissioning development.
                <ul>
                  <li id="sec-2.4.1.1"><span class="todo TODO"> TODO</span>
                    ATLiS to test optics measurements at 1E03/2S10
                    (before Jan 2011 optics) <br>
                  </li>
                  <li id="sec-2.4.1.2"><span class="todo TODO"> TODO</span>
                    Work up detailed tabular strawman, cost/resource
                    estimates (Benesch) <br>
                  </li>
                  <li id="sec-2.4.1.3"><span class="todo TODO"> TODO</span>
                    Cross-check injector harp optics match measurement
                    consistency <br>
                  </li>
                </ul>
              </li>
            </ul>
          </div>
        </div>
        <div id="outline-container-2.5" class="outline-3">
          <h3 id="sec-2.5"><span class="section-number-3">2.5</span>
            Anything else? </h3>
          <div class="outline-text-3" id="text-2.5">
            <ul>
              <li id="sec-2.5.1">Optics: Hofler, Kazimi <br>
              </li>
            </ul>
          </div>
        </div>
      </div>
      <div id="postamble">
        <p class="date"> Date: 2010-12-13 15:37:41 EST</p>
        <p class="creator">HTML generated by org-mode 6.33x in emacs 23</p>
      </div>
    </div>
    <meta http-equiv="content-type" content="text/html;
      charset=ISO-8859-1">
    <br>
    <pre class="moz-signature" cols="72">-- 
Director of Accelerator Operations
Jefferson Lab           Office: (757) 269-6268  
12000 Jefferson Avenue  Cell:   (757) 876-6289  
Newport News, VA 23606  Email:  <a class="moz-txt-link-abbreviated" href="mailto:freyberg@jlab.org">freyberg@jlab.org</a>
</pre>
  </body>
</html>